# A.I.R.: Adaptive, Iterative, and Reasoning-based Frame Selection For Video Question Answering

### v1

@@ A.4.4 Additional Analysis of Method Ablations
Influence of Method Components. To validate the contribution of each component within the A.I.R. framework, we conduct a systematic ablation study, with results presented in Tab. 4. Our full method (#2) achieves 68.2% accuracy, a significant improvement of +2.6% over the Uniform Sampling baseline (#1), while also using fewer frames on average. Disabling individual components generally degrades performance, confirming their synergistic importance. The most critical component is the Reasoning-based VLM Analysis (#7); its removal causes the largest accuracy drop to 66.0% and increases the frame count to 32. The Interval Potential Ranking (#6) and Adaptive Initial Sampling (#5) are also highly impactful, with their removal reducing accuracy to 66.7% and 66.9%, respectively. Notably, replacing our adaptive budget with a fixed 32-frame budget (#3) yields a marginal 0.1% increase in accuracy to 68.3%. However, we selected our adaptive approach as the default because it strikes a superior efficiency-performance trade-off: it achieves nearly identical top-tier performance while using 22.5% fewer frames on average (24.8 vs. 32.0), making it a more practical and scalable solution.
@@
A.4.5 Additional Efficiency Analysis and Hyperparameter Ablations
@@
Max Sampling
Budget ($V_{\max}$)
Candidate Pool
Size ($C$)
Max
Iterations ($\mathcal{I}_{\max}$)
Time (s)
Accuracy

### v2

@@ A.4.4 Additional Analysis of Method Ablations
Influence of Method Components. To validate the contribution of each component within the A.I.R. framework, we conduct a systematic ablation study, with results presented in Tab. 4. Our full method (#2) achieves 68.2% accuracy, a significant improvement of +2.6% over the Uniform Sampling baseline (#1), while also using fewer frames on average. Disabling individual components generally degrades performance, confirming their synergistic importance. The most critical component is the entire Iterative Frame Selection stage (#6); its removal causes the largest accuracy drop to 65.2%, demonstrating that the iterative refinement process is essential for identifying relevant frames. The Reasoning-based VLM Analysis (#8) is also highly impactful, with its removal reducing accuracy to 66.0% and increasing the frame count to 32. The Interval Potential Ranking (#7) and Adaptive Initial Sampling (#5) are similarly important, with their removal reducing accuracy to 66.7% and 66.9%, respectively. Notably, replacing our adaptive budget with a fixed 32-frame budget (#3) yields a marginal 0.1% increase in accuracy to 68.3%. However, we selected our adaptive approach as the default because it strikes a superior efficiency-performance trade-off: it achieves nearly identical top-tier performance while using 22.5% fewer frames on average (24.8 vs. 32.0), making it a more practical and scalable solution.
@@
A.4.5 Additional Efficiency and Hyperparameter Analysis
@@
# Max Sampling Budget ($V_{\max}$) Candidate Pool Size ($C$) Max Iterations ($\mathcal{I}_{\max}$) Time (s) Accuracy (%)
@@
A.4.6 Comprehensive Hyperparameter Ablations
To validate the robustness of A.I.R. across different hyperparameter configurations, we conduct a systematic ablation study on LongVideoBench (LVB) and NextQA using InternVL3-8B. As shown in Tab. 14, we independently vary each of the five key hyperparameters while keeping others fixed at their default values: $\gamma=0.7$ (GMM coefficient), $\alpha=15$ (minimum frame distance), $\beta=1.5$ (LDS growth factor), $l_{\min}=20$ (minimum event length for pruning), and $d_{\min}=2.0$s (minimum gap for merging events).
Individual Parameter Analysis. Our results demonstrate strong robustness across all hyperparameters. For the GMM coefficient $\gamma$ (rows 1-3), we observe that moderate values (0.5-0.7) yield the best performance, with $\gamma=0.5$ and $\gamma=0.7$ both achieving 82.6% on NextQA. Extreme values ($\gamma=0.3$ or $\gamma=0.9$) lead to slightly degraded performance (-0.3% to -0.4%), as they either select too many low-relevance frames or miss important events. For minimum frame distance $\alpha$ (rows 4-6), larger values (20-25) prevent redundant sampling but may miss fine-grained details, while smaller values (10) enable denser coverage. The LDS growth factor $\beta$ (rows 7-9) shows consistent performance across the tested range (1.2-2.0), with moderate values (1.5-1.7) performing slightly better. The pruning threshold $l_{\min}$ (rows 10-12) and merging gap $d_{\min}$ (rows 13-14) both demonstrate stability, with performance varying by less than 1.0% across different settings.
Strategy-Based Combinations. We further test four strategy combinations (rows 15-18) designed for different video characteristics: (15) Short-Video strategy with dense sampling ($\gamma=0.5$, $\alpha=10$, $l_{\min}=15$, $d_{\min}=1$); (16) Long-Video strategy with sparse sampling ($\alpha=25$, $l_{\min}=30$, $d_{\min}=3$); (17) Conservative strategy with strict thresholds ($\gamma=0.3$, $\alpha=20$); and (18) Aggressive strategy with loose thresholds ($\gamma=0.9$, $\alpha=20$). All combinations maintain competitive performance (61.7-62.9% on LVB, 81.8-82.7% on NextQA), demonstrating that A.I.R. can be flexibly adapted to different video types while preserving effectiveness.
Robustness Analysis. Critically, all 18 configurations maintain performance within $\pm$1.1% of the default setting on both benchmarks, confirming that A.I.R. is highly robust to hyperparameter choices.
Table 14: Systematic ablation study of A.I.R. hyperparameters using InternVL3-8B. Default configuration (gray) uses parameters from Appendix A.3.1: $\gamma=0.7$, $\alpha=15$, $\beta=1.5$, $l_{\min}$=20, $d_{\min}$=2.0s. We vary each parameter independently to isolate effects. All configurations maintain performance within ±1.1% of baseline, demonstrating strong robustness. Rows 15-18 test strategy combinations: Short-Video (dense sampling), Long-Video (sparse sampling), Conservative (strict), Aggressive (loose). All use $V_{\max}=32$, $C=12$, $\mathcal{I}_{\max}=6$.
   #    $\gamma$    $\alpha$    $\beta$    $l_{\min}$    $d_{\min}$    LVB    NextQA
[…]
   $\gamma$ ablation (GMM coefficient)
   1    0.3    15    1.5    20    2    62.2    82.3
   2    0.5    15    1.5    20    2    62.6    82.6
   3    0.9    15    1.5    20    2    62.0    82.2
   $\alpha$ ablation (min_frame_distance)
   4    0.7    10    1.5    20    2    62.4    82.8
   5    0.7    20    1.5    20    2    62.8    82.1
   6    0.7    25    1.5    20    2    62.9    82.0
   $\beta$ ablation (LDS growth factor)
   7    0.7    15    1.2    20    2    62.4    82.4
   8    0.7    15    1.7    20    2    62.6    82.5
   9    0.7    15    2.0    20    2    62.1    82.1
   $l_{\min}$ ablation (Pruning)
   10    0.7    15    1.5    15    2    62.0    82.4
   11    0.7    15    1.5    25    2    62.8    82.1
   12    0.7    15    1.5    30    2    62.7    81.8
   $d_{\min}$ ablation (Merging)
   13    0.7    15    1.5    20    1    62.3    82.8
   14    0.7    15    1.5    20    3    62.6    82.3
   Strategy-based combinations
   15    0.5    10    1.5    15    1    61.7    82.7
   16    0.7    25    1.5    30    3    62.9    81.8
   17    0.3    20    1.7    20    2    61.8    82.0
   18    0.9    20    1.2    20    2    62.1    81.9